what is the maximum product of two numbers whose sum is -10?
Answer by stanbon(75887)

what is the maximum product of two numbers whose sum is -10?
x + y = -10
------
Product = x*y = x(-10-x) = -10x - x^2
-------
Max occurs when x = -b/(2a) = 10/(2*-1) = -5
------
Then y = -10-(-5) = -5
-----
Ans: x = y = -5
----------------------
Cheers,
Stan H.
---------------------